OpenStudy (anonymous):

WILL FAN AND MEDAL!!! solve the system of equations 2x+3y=1 y=3x+15

OpenStudy (anonymous):

@meganmerie

OpenStudy (anonymous):

PLS help @texaschic101 and @fallenangelorchid

OpenStudy (fallenangelorchid):

@KamiBug

OpenStudy (anonymous):

lol

OpenStudy (kamibug):

In the second equation, y is already isolated, so we know it's value. That means we can plug in 3x+15 for y in the first equation to solve for x. Once we know what x is, we can substitute it into the second equation to find y. :) Substitute. 2x+3(3x+15)=1 Distribute. 2x+9x+45=1 Combine Like Terms. 11x = -44 Divide both sides by 11. x = -4 <--- Now substitute -4 for x in the second equation to find y. y=3(-4)+15 y = -12 + 15 y = ? <---

OpenStudy (anonymous):

so how are you going to find the two ordered pairs

OpenStudy (kamibug):

x = -4 and y = 3 That means the solution is at (-4, 3). :)

OpenStudy (anonymous):

thank you @KamiBug